write an equation of the line containing the given poin tand perpendicular to the given line. Express your answer in the form y=mx+b
(5,6);2x+y=7

Answer by mananth(16946)

2 x + 1 y = 7
Find the slope of this line

1 y = -2 x + 7
Divide by 1
y = -2 x + 7
Compare this equation with y=mx+b
slope m = -2.00

The slope of a line perpendicular to the above line will be the negative reciprocal 0.5
m1*m2=-1
The slope of the required line will be 1/2

m= 1/2 ,point ( 5 , 6 )
Find b by plugging the values of m & the point in
y=mx+b
6 = 5/2 + b
b= 7/2
m= 1/2
The required equation is y = 1/ 2 x + 7/ 2
m.ananth@hotmail.ca